HUBBLE SPACE TELESCOPE DAILY REPORT #830 PERIOD COVERED: 0800 EST Daily Status Report as of 055/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FOS/BL 3545 (Evolution of Ultraviolet Spectra in Early Type Galaxies) The Faint Object Spectrograph was used to make four spectral observations of early type galaxy 55W037. It observations was made in parallel with the prime science observation. All four observations were executed in course track. They were received and displayed in the OSS. They were all reported to look good. 1.2 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Both the wide Field Camera and the Planetary Camera were used to make internal flat observations to look for new forms of contamination (i.e., measles and daisies), and the return of quantum efficiency hysteresis. Both images were received and examined in the OSS. They were reported to look normal. 1.3 Completed Two Sets of PC 4714 (WF/PC Calibration: PC Internal Delta Flats) The Planetary Camera was used to make a total of three internal delta flat images. These images will be used for the restoration of Earth flats, which were are altered due to changes in quantum efficiency which occur after WFPC decontamination operations. All three images were displayed, and reported to look okay. 1.4 Completed WFC 4713 (WF/PC Calibration: Internal Delta Flats) The Wide Field Camera was used to take one internal flat observation. The image was displayed in the OSS, and was reported to look okay. 1.5 Completed PC 3769 (Spectra of Inner Coma of Comet Chrion) The Planetary Camera was used to make 13 observations of the comet/object Chiron to study the inner coma region, and dust near the edge of its exopause. The first nine observations were received and displayed. They were reported to look okay. The next four observations were executed in course track. These were point spread function observations and the engineering telemetry rate was only 4K for these observations. The observations were reported to look okay, however, they were filed as being bad because the engineering telemetry was in the low data rate. 1.6 Completed GHRS/2 3862 (Excitation of Atmospheres of Planetary Satellites) The High Resolution Spectrograph was used to make eight spectral observations of Jupiter's moon Io while it entered the planet's shadow. All eight observations were received and where possible displayed. The onboard target acquisition was successful. The spectrum was reported to look normal. 1.7 Completed Two Sets of FOS/RD 4784 (Transition Comets: UV Search for OH) The Faint Object Spectrograph was used to make two observations of the asteroid HIDALGO looking for OH emission lines at 3085 A0. All three observations were received and displayed. The first observation was the onboard target acquisition. It was successful. The second was the ultraviolet spectral observation. It was reported to look normal. 1.8 Completed PC 3230 (PC PSF and Focus Monitoring 2) The Planetary Camera was used to make point spread function observations and focus change observations using standard star SAO62896. These changes may have been caused by desorption or other changes in the WFPC system. The observations were executed as scheduled, and no problems were reported. 1.9 Initiated HSP/UV1 4209 (HSP Detector Sensitivity Test) Observations were initiated using the High Speed Photometer to measure the sensitivity of the four HSP Image Detector Tubes. The calibration source is standard star VID998. The observations are currently under way, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Note: A special FGS #2 analysis and review meeting to study the star servo A anomaly was held yesterday at GSFC (Tuesday, Feb. 23,1993). This meeting lasted all day and included experts from both government and industry on bearings, motors, and servos. The meeting will continue this afternoon (Wednesday, Feb. 24) at 1 p.m. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of HSP/UV1 4209 (HSP Detector Sensitivity Test) 3.2 Complete Three Sets of WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) 3.3 Complete HSP/UV2 1095 (Variability of High Luminosity Stars) 3.4 Complete FGS 4064 (Crit. Test of Galactic Escape Velocities at R(Sun)) 3.5 Complete Two Sets of WFC 4105 (HST Medium Deep Survey: High Latitude) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete PC 3209 (PC Images of Eta Carnae Core) 3.8 Complete PC 3603 (Snapshots of Protoplanetary Nebulae) 3.9 Complete FOS/RD 4495 (FOS GIMP Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-